OpenStudy (anonymous):

What are the solutions of the equation? z² – 6z – 27 = 0

OpenStudy (anonymous):

(3,9) (3,-9) (-3,9) (-3,-9)

OpenStudy (cyber405):

First find the delta: b^2-4ac then for finding the answer use this....\[(-b \pm \sqrt{\Delta})\div2a\]

OpenStudy (anonymous):

huh?

OpenStudy (cyber405):

let me show the steps...

OpenStudy (anonymous):

yes please!

OpenStudy (cyber405):

You can write this equation like this: \[az ^{2}+bz+c\] ok?

OpenStudy (anonymous):

yeah

OpenStudy (cyber405):

For finding answers of this equation you will need Delta... \[\Delta=b ^{2}-4ac\] if delta>0 then you have two answers... if delta=0 then you have one answer... if delta <0 then you don't have answer for the equation...

OpenStudy (cyber405):

For finding answer you should use this: \[(-b \pm \sqrt{\Delta})\div2a\]

OpenStudy (anonymous):

i have no clue how. that is gibberish to me

OpenStudy (cyber405):

in this equation a=1 and b= -6 and c= -27 ok?

OpenStudy (anonymous):

yeah i got that part

OpenStudy (cyber405):

Delta = \[(-6)^{2}-4(1)(-27)=144\] ok?

OpenStudy (anonymous):

nope lost me

OpenStudy (cyber405):

Δ=\[b ^{2}-4ac\] here b=-6 and a=1 and c=-27 ....just put the nimbers in the delta equation...ok?

OpenStudy (anonymous):

ok so what do i do with the 144?

OpenStudy (cyber405):

delta is 144 and we need it to find the answers... Delta > 0 then we should have two answers... Answers = \[(-b+\sqrt{\Delta})\div2a\] and \[(-b-\sqrt{\Delta})\div2a\] Just put numbers and get the answers... answers are....???

OpenStudy (cyber405):

-3 and 9 got them?

OpenStudy (anonymous):

no clue how you got them but thanks!

OpenStudy (cyber405):

Look i will show you... \[(-(-6)+\sqrt{144})\div2(1)=9\] \[(-(-6) -\sqrt{144})\div2(1)\]=-3 just remember b=(-6) , a=1 , c=-27 , delta=144

OpenStudy (cyber405):

Need more explain...?